AGw3 Calibration at LFBG with LUCI1: 10 April 2012 UT

Observer: JHill (LBTO), and eventually DThompson (Tucson), DMiller (Tucson)
Telescope Operator: GBechetti (LBTO)
Software Support: (Tucson)
Telescope Support: JUrban (LBTO)
Instrument Support: DOfficer (LBTO)


After some fits and starts, we got the hotspot and guide probe realigned for AGw3. Handed over the telescope to LBTI at midnight, as LUCI1 is not ready to take transformation data.

LBTI had some struggles with pointing of telescope optics and acquiring the images. See Issues 4035 (crazy collimation) and 4036 (dx.terc).


LUCI1 has been remounted on the telescope, and the telescope has been rebalanced.

01:57 Doug O. turns on the LUCI1 ROE switch.



01:50 Open to partly cloudy sky. T=8 degC D=-12 degC, Wind light from West.

01:55 At zenith for twilight work

M3 Alignment confirmation at zenith

setGSsfp left

readGuideCam left -e 300

Images are in /OldRepository/AGW_Data/20120410/

wfs 0004 360 deg
wfs 0005 270 deg
wfs 0006 180 deg
wfs 0007 90 deg

The outer edge of the pupil doesn't seem very well defined against the twilight sky. However, based on the central shadow of the pupil the chief ray coming from the tertiary looks decently well aligned. This is as we expect since nothing has changed on the tertiary. But, based on our preset after this, SX M3 might not have been in position.

Pointing & Rough Collimation

We are authorized as LUCI on Left and LBTI on the right in order to test out binocular collimation updating.

02:14 ACQUIRE preset - seems to have hung on collimating M3.

02:20 Restart OSS and Init SX M3. Had to $$$ reset UMAC to get it to Init.

02:25 Now the preset works. ptspiral finds the star with lots of defocus and coma. Why?

02:42 Switch back to LUCI and None authorization since Steve's pointing tool can't yet handle mixed mode pointing models. The pointing offsets in PSF did not clear when we did this. (That must be a 2012D feature.)

02:45 ACQUIRE preset to SX - but still some problems finding the star, so something else must be wrong. Perhaps we are too close to zenith at EL=88 and rising.

02:58 ACQUIRE preset -- Steve's pointing tool doesn't seem to work for LUCI either. Geno has to use the command line to adjust pointing.

03:10 Oops..... The Retro guys had disabled the collimation LUT on M2. That was the cause of our pointing, collimation, focus issues.

03:11 ACQUIRE preset - Now the star is near-by, less far from focus and not comatic.

03:18 ACTIVE preset to an off-axis guide star on this same target (2nd mag M5 star) - but stopping WFS manually.

Doing eyeball active optics to get the image smaller than the pinhole.

Rough Guiding Hotspot Adjustment

03:52 guidecam_hotspot_y changed from 290 to 220. This is close, but not yet aligned at the sub-arcsec level.

03:53 stop and restart GCSL to be sure the change takes hold

03:54 resend preset, but clouds have thinned and guidestar is too bright for WFSing. (needed 3 sec exposures)

03:59 We didn't find the guide star for the BS star.

04:04 ACQUIRE preset to adjust pointing

04:08 ACTIVE preset on WT10 star, but couldn't see the 13th mag guide star - maybe more clouds are rolling through.

04:16 ACQUIRE preset -- pointing was good

04:18 ACTIVE misses the guide star - why?

04:20 ACQUIRE preset -- then rotate the rotator 360 deg. This makes the star follow a circle more than 10 arcsec in diameter. That is why we were having troubles finding guidestars at arbitrary position angles.

Measuring Center of Rotation and Probe Adjustment

04:26 ACQUIRE preset to BS star on-axis.

stopAcquire left (stops the movie acquisition)

readGuideCam left 80000 (while rotator moving 7-367 deg) , but that was taking 80000 images...... thru guider 0017

probe is at native x=-0.381 y=611.48

readGuideCam left -e 80000 guider 0019

Approximate center of the circle is x=133, y=18 on the binned x2 CCD. I think this means we have a problem with the alignment.

Move probe to -0.381 606.0 mm

04:58 readGuideCam left -e 90000 guider 0021

And the bad news is that this moved the center of rotation down off of the guider chip, and not up to binned y=110 where we wanted it.

Move probe to -0.381 612.5

05:08 readGuideCam left -e 90000 guider 0022 Center is at x=131 y=26 in binned x2 pixels.

Move probe to -0.381 615.5

05:15 readGuideCam left -e 90000 guider 0023 Center at x=130 y=78 in binned x2 pixels.

Move probe to -0.381 617.3

05:23 readGuideCam left -e 92000 guider 0024 Center at x=132.6 y=109 in binned x2 pixels.

Guide probe position x: -25.000000 y: 1030012.000000 x: -0.370000 y: 617.299000 (These are the encoder coords to center it.)

Guide probe position x: -25.000000 y: 1005062.000000 x: -0.367000 y: 612.501000 (These are the xy coords.)

05:45 edit oacontrol.conf
617.3 mm - 612.5 mm = 4800 microns (the number to subtract from ooffset)
Change ooffset from 418220 to 413420 microns. Change y_guide from 289.3 to 220.0.

05:52 stopGCS, stopAGW, restart oacontrol, startAGW, start GCS, resend preset

(see also the previous alignment activities on 20120204 and 20111101.)

06:02 ACQUIRE preset to recenter.

Collimation and Pupil Position

06:06 ACTIVE preset on-axis -- image improves to 0.4 arcsec FWHM on guider.

Good on-axis WFS images are at 06:10 UT +- 1 minute.

06:11 ACTIVE preset off-axis (on the y-axis near the edge of the field x=-1.2 y=432.5 mm)

06:16 Guiding runs away because of bright lights on guider. Sextractor picked the bright gradient in lower right corner. Moon? Tomato Farm? Also lots of light on the WFS outside M2. Close vent doors. and the stray light is gone. This is Issue 4038.

06:19 resend preset.

06:22 send preset with PA=268 (248 was on the y-axis)

guidecam_hotspot_x      250.0           double  # ccd x coordinate of pixel that represents center of light in WFS pupil 289.0 (20120410 JMH)
guidecam_hotspot_y      220.0           double  # ccd y coordinate of pixel that represents center of light in WFS pupil 252.0 (20120410 JMH)
WFScam_hotspot_x        385.1           double  # ccd x coordinate of pixel that represents leftmost x-center of WFS spots (20120204 DLM)
WFScam_hotspot_y        235.0           double  # ccd y coordinate of pixel that represents leftmost y-center of WFS spots (20120410 JMH)

The guide hotspot position was also updated in oacontrol.conf in both the x_0,y_0 and the x_guide,y_guide variables.

Precise Guiding Hotspot Adjustment

06:29 ACTIVE preset off-axis. Changing fields to get a faint guide star to make precise hotspot alignment.

06:41 centroids where the hotspot should be 30.8 37.9

06:43 centroids recentered at the old hotspot 38.4 38.6

06:44 Adjust guidecam_hotspot_x from 265 to 250 and restart GCSL. Note that we did not attempt to make a theta adjustment to the probe for this last small refinement in X.

06:48 resend ACTIVE preset

Transformation Data

LUCI1 isn't checked out yet, so we can't take transformation data tonight.

06:50 Handover telescope to LBTI (first for their two hours that are scheduled for tonight, and later hoping that they'll pay us back 4 hours tomorrow)

09:09 Installed Doug's measured on-axis pupil position in LUCI_L.cfg --- WFScam_hotspot_y changed from 247 to 235 .

-- JohnHill - 10 Apr 2012
Topic revision: r7 - 19 Apr 2012, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback